OpenStudy (anonymous):

**Calc 2 Help Needed** Solve the initial value problem below: (t^2 + 2t)dx/dt = 2x + 5 with t, x>0 and x(1) = 1

OpenStudy (anonymous):

Didn't i solved this

OpenStudy (anonymous):

I needed additional assistance.. I understand to split and integrate, but don't know where to head afterwards?

OpenStudy (anonymous):

Once you integrate you get \[\ln{(2x+5)}=\frac{1}{2}\ln(\frac{t}{t+2})+c\] put x=1 and t=1 to find c

OpenStudy (anonymous):

\[c=\ln(7\sqrt3)\]

OpenStudy (anonymous):

ok makes sense so far..

OpenStudy (anonymous):

That is it nothing is unknown now

OpenStudy (anonymous):

so how do I slove for x(t) now?

OpenStudy (anonymous):

ln(7) - (ln(1/3))/2??

OpenStudy (anonymous):

ahh that didn't work online.

OpenStudy (anonymous):

are you still there??